OpenStudy (anonymous):

4) The DIAMETER of a sphere is doubled. What is the new volume of the sphere? A) 83πr3 B) 163πr3 C) 323πr3 D) 643

OpenStudy (amistre64):

what is the usual formula for the volume of a sphere?

OpenStudy (anonymous):

We need the original volume or diameter.

OpenStudy (anonymous):

thats what im trying to see

OpenStudy (amistre64):

dont need the original stuff; just let r = d/2 in the usual formula then let d = 4d

OpenStudy (amistre64):

pfft, d = 2d :)

OpenStudy (amistre64):

so, since r = d/2, and twice d is 2d the new r is simply, r = 2d/2 = d = (2r)

OpenStudy (amistre64):

i recall the r being cubed in the usual format soo, (2r)^3 = ?

OpenStudy (anonymous):

the answer is b

OpenStudy (amistre64):

i do not believe it is b, what is your reasoning?

OpenStudy (anonymous):

sa=4pi r2 =4pi(2r)^2 4pi time 4r2 16pi r2

OpenStudy (amistre64):

volume is not surface area

OpenStudy (anonymous):

its easy we lookin for the surface area

OpenStudy (anonymous):

not volume

OpenStudy (amistre64):

your surface area equation is rather nice tho :) but we should do that for volume of a sphere ... then your post was incorrect when it asked for volume?

OpenStudy (amistre64):

all you options have an r^3 in them as well